# A Szegö limit theorem for a class of Toeplitz operators on the Bergman space of the unit ball with singular symbols Daniel Ivan Ramirez Montaño

### v1

@@ 6.2 Schatten norm estimates
being finite. For $p\geq 1$, $\left|\left|\cdot\right|\right|_{\scriptscriptstyle S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))}$ is a norm in $S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))$, while when $10<p<$ 1 it is a quasi-norm satisfying
$\left|\left|A+B\right|\right|_{\scriptscriptstyle S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))}^{p}\leq 2\left(\left|\left|A\right|\right|_{\scriptscriptstyle S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))}^{p}+\left|\left|B\right|\right|_{\scriptscriptstyle S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))}^{p}\right)\ .$
Let $a\in C^{2}_{0}(\Gamma)$. If $\widehat{T}_{a\mathrm{d}\sigma}$ is a Toeplitz operator, it can be expressed as the sum of three Toeplitz operators with positive symbols (see the beginning of the proof of 5.2). Then, as $\left|\left|\cdot\right|\right|_{\scriptscriptstyle p}$ is at least a quasi-norm, we will have by Lemma 5.9 that
$\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}((\widehat{T}_{a\mathrm{d}\sigma}^{\ast}\widehat{T}_{a\mathrm{d}\sigma})^{\frac{p}{2}})$
is bounded when $\alpha\rightarrow\infty$.
Lemma 6.2.
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{2}_{0}(\Gamma)$. For any $p>0$,
$\lim\limits_{\alpha\rightarrow\infty}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}\big{(}({\widehat{T}_{a\mathrm{d}\sigma}}^{\ast}\widehat{T}_{a\mathrm{d}\sigma})^{\frac{p}{2}}\big{)}=\frac{1}{(2p)^{\frac{d^{\prime}}{2}}}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{p}\ \mathrm{d}\sigma\ .$
Proof.
For any Toeplitz operator, ${\widehat{T}_{a\mathrm{d}\sigma}}^{\ast}=\widehat{T}_{\overline{a}\mathrm{d}\sigma}$ and so we have for $m\in\mathbb{N}$
$\mathrm{Tr}(({\widehat{T}_{a\mathrm{d}\sigma}}^{\ast}\widehat{T}_{a\mathrm{d}\sigma})^{m})$ $=\frac{\left(\frac{\alpha}{\pi}\right)^{\frac{d}{2}}}{(4m)^{\frac{d^{\prime}}{2}}}\int_{\Gamma}\frac{|a(\xi)|^{2m}}{(1-|\xi|^{2})^{2m(n+1)}}\ \mathrm{d}\sigma(\xi)\ \left(1+O\left(\frac{1}{\alpha}\right)\right)$
$=\frac{\left(\frac{\alpha}{\pi}\right)^{\frac{d}{2}}}{2^{d^{\prime}}}$ $\int_{\Gamma}\mathcal{Q}_{\frac{d^{\prime}}{2}}(s^{m})\left(\frac{|a(\xi)|^{2}}{(1-|\xi|^{2})^{2(n+1)}}\right)\mathrm{d}\sigma(\xi)\left(1+O\left(\frac{1}{\alpha}\right)\right)\ .$
Make $A:={\widehat{T}_{a\mathrm{d}\sigma}}^{\ast}\widehat{T}_{a\mathrm{d}\sigma}$. Using the same arguments as in Corollary 5.2 and the discussion at the beginning of Subsection 5.2 we can find a constant $R$ such that $\sigma(A)\subseteq[0,R^{2}]$ for all $\alpha>0$.
By the linearity of $\mathcal{Q}_{\frac{d^{\prime}}{2}}$ we get
$\lim\limits_{\alpha\rightarrow\infty}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}(P(A))=\frac{1}{2^{d^{\prime}}}\int_{\Gamma}\mathcal{Q}_{\frac{d^{\prime}}{2}}(P)\left(\frac{|a(\xi)|^{2}}{(1-|\xi|^{2})^{2(n+1)}}\right)\ \mathrm{d}\sigma(\xi)$ (41)
for any $P$ polynomial without constant term. Let $\epsilon>0$. We compute
$\left|\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}\big{(}A^{\frac{p}{2}}\big{)}-\frac{1}{(2p)^{\frac{d^{\prime}}{2}}}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{p}\mathrm{d}\sigma\right|\leq\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\bigg{|}\mathrm{Tr}\big{(}A^{\frac{p}{2}}-P(A)\big{)}\bigg{|}$
$\qquad+\left|\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}\big{(}P(A)\big{)}-\frac{1}{2^{d^{\prime}}}\int_{\Gamma}\mathcal{Q}_{\frac{d^{\prime}}{2}}(P)\left(\frac{|a(\xi)|^{2}}{(1-|\xi|^{2})^{2(n+1)}}\right)\mathrm{d}\sigma\right|$
$\qquad\qquad+\frac{1}{2^{d^{\prime}}}\left|\int_{\Gamma}\mathcal{Q}_{\frac{d^{\prime}}{2}}(s^{\frac{p}{2}}-P)\left(\frac{|a(\xi)|^{2}}{(1-|\xi|^{2})^{2(n+1)}}\right)\mathrm{d}\sigma\right|=:\mathring{I}_{7}+\mathring{I}_{8}+\mathring{I}_{9}$ (42)
for any such polynomial. We will consider a $P$ of the form $P=P_{1}P_{2}$ with $P_{1}$ and $P_{2}$ two other polynomials without constant term. By doing this we have
$|s^{\frac{p}{2}}-P|\leq|s^{\frac{p}{4}}||s^{\frac{p}{4}}-P_{1}(s)|+|P_{1}(s)||s^{\frac{p}{4}}-P_{2}(s)|\ .$
Using this, we estimate $\mathring{I}_{7}$ by
$\mathring{I}_{7}$ $\leq\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\bigg{|}\mathrm{Tr}\big{(}A^{\frac{p}{4}}(A^{\frac{p}{4}}-P_{1}(A))\big{)}\bigg{|}+\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\bigg{|}\mathrm{Tr}\big{(}P_{1}(A)(A^{\frac{p}{4}}-P_{2}(A))\big{)}\bigg{|}$
$\leq\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}(A^{\frac{p}{4}})\big{|}\big{|}s^{\frac{p}{4}}-P_{1}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}+\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}(P_{1}(A))\big{|}\big{|}s^{\frac{p}{4}}-P_{2}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]},$
and, applying c) of Proposition 5.8, $\mathring{I}_{9}$ by
$\mathring{I}_{9}$$\leq$ $\frac{1}{2^{\frac{d^{\prime}}{2}}\left(\frac{p}{4}\right)^{\frac{d^{\prime}}{2}}}\left|\left|\frac{s^{\frac{p}{2}}-P}{s^{\frac{p}{4}}}\right|\right|_{\scriptscriptstyle L^{\infty}[0,R^{2}]}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{\frac{p}{2}}\mathrm{d}\sigma(\xi)$
$\leq$ $\left(\frac{2}{p}\right)^{\frac{d^{\prime}}{2}}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{\frac{p}{2}}\mathrm{d}\sigma(\xi)$
$\qquad\cdot\left(\big{|}\big{|}s^{\frac{p}{4}}-P_{1}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}+\big{|}\big{|}P_{1}s^{-\frac{p}{4}}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}\big{|}\big{|}s^{\frac{p}{4}}-P_{2}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}\right)$
for all $\alpha>0$. We can now choose $P_{1}$ such that
$\big{|}\big{|}s^{\frac{p}{4}}-P_{1}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}<\frac{\epsilon}{6}\min\left\{{\scriptstyle\left(\sup\limits_{\alpha>0}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}(A^{\frac{p}{4}})\right)^{-1},\ \Big{(}\left(\frac{2}{p}\right)^{\frac{d^{\prime}}{2}}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{\frac{p}{2}}\mathrm{d}\sigma(\xi)\Big{)}^{-1}}\right\}$
and $P_{2}$ such that
$\big{|}\big{|}s^{\frac{p}{4}}-P_{2}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}<\frac{\epsilon}{6}\min\bigg{\{}{\scriptstyle\left(\sup\limits_{\alpha>0}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}(P_{1}(A))\right)^{-1}},\\ {\scriptstyle\Big{(}\left(\frac{2}{p}\right)^{\frac{d^{\prime}}{2}}\big{|}\big{|}s^{-\frac{p}{4}}P_{1}\big{|}\big{|}_{\scriptscriptstyle L^{\infty}[0,R^{2}]}\int_{\Gamma}\left(\frac{|a(\xi)|}{(1-|\xi|^{2})^{n+1}}\right)^{\frac{p}{2}}\mathrm{d}\sigma(\xi)\Big{)}^{-1}}\bigg{\}}\ .$
This ensures that $\mathring{I}_{7}<\frac{\epsilon}{3}$ and $\mathring{I}_{9}<\frac{\epsilon}{3}$. Finally for these fixed $P_{1}$ and $P_{2}$ we know from (41) that we can take $\alpha$ large enough so $\mathring{I}_{8}<\frac{\epsilon}{3}$, which together with (6.2) concludes the proof. ∎
The desired result describing the Schatten norms follows as a consequence.
@@
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{2}_{0}(\Gamma)$. For any $p>0$,
@@ 6.3 An example
For $r\in(0,1)$ let $\Gamma=r\mathbb{S}^{1}=\{z\in\mathbb{C}:|z|=r\}$. If we use the parametrization $\gamma:(0,1)\rightarrow\Gamma$, $\gamma(\theta)=re^{2\pi i\theta}$, then $\mathrm{d}\sigma=2\pi\frac{r}{1-r^{2}}\mathrm{d}\theta$. Notice that in this case $d=1$ and $n=1$, thus $d^{\prime}=1$. Let $a\equiv 1$ in $\Gamma$ and $T_{r}:=T_{1\mathrm{d}\sigma}$ be the associated Toeplitz operator. Then
@@ 6.3.1 Operator norm
from the known asymptotic formula $\mathrm{\mathbf{B}}\left(x,y\right)\sim\sqrt{2\pi}\frac{x^{x-\frac{1}{2}}y^{y-\frac{1}{2}}}{(x+y)^{x+y-\frac{1}{2}}}$ for large $x$ and $y$, which is in turn an application of Stirling’s formula. Substituting this in the expression for $\lambda_{m^{*}}$ in (43) we get
@@ 6.3.2 Eigenvalue distribution comparison
Keeping with $\widehat{T}_{r}$ as defined above, given an interval $[t_{1},t_{2}]\subset\left(0,\frac{1}{(1-r^{2})^{2}}\right)$ Corollary 6.1 simplifies to
@@
To test this result, for fixed values of $r\in(0,1)$ and $t_{1}<t_{2}\leq\frac{1}{(1-r^{2})^{2}}$, and distinct values of $\alpha$, we have numerically counted the number of eigenvalues of $\widehat{T}_{r}$ in $[t_{1},t_{2}]$ using (43). For a pair of selected values, the comparisons between these numerical outputs and the expected results given by (44) and (45) are presented in the next two tables.
@@ B.1 The determinant as a block matrix
Recall from (23) that computing the determinant of the Hessian of $\varphi(t,\boldsymbol{0})$ relies on computing that of $\mathcal{O}_{m-1}$, which we have defined as the $(m-1)d\times(m-1)d$ matrix that expressed as a $(m-1)\times(m-1)$ block matrix with blocks of size $d\times d$ has the form
$\mathcal{O}_{m-1}=\begin{pmatrix}2I&-I-iW&{\scriptstyle\cdots}&0\\ -I+iW&2I&{\scriptstyle\cdots}&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&{\scriptstyle\cdots}&2I&-I-iW\\ 0&{\scriptstyle\cdots}&-I+iW&2I\end{pmatrix},$
where $W:=G^{-1}H$.
Let $\mathcal{R}$ stand for the ring of $d\times d$ complex matrices generated by $I$ and $W$. Then $\mathcal{O}_{m-1}$ can also be regarded as a $(m-1)\times(m-1)$ matrix with entries in $\mathcal{R}$. As $\mathcal{R}$ is commutative, any such matrix possesses a well-defined determinant in $\mathcal{R}$, which we will denote $\mathrm{det}_{\mathcal{R}}$, and that shares all the usual properties with determinants of square matrices with complex entries. The relation between this determinant and the usual one $\det\mathcal{O}_{m-1}$ is given in Theorem 1 in [12]. From that theorem it follows that
@@
Notice that the right-hand side of the equation makes sense since $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{m-1}$ is an element of $\mathcal{R}$, and thus a $d\times d$ matrix with complex entries. We will use this fact to compute $\det\mathcal{O}_{m-1}$. This demonstration is the same as the one in [11].
@@
Let $q\in\mathbb{N}$. Then
$\mathrm{det}_{\mathcal{R}}\mathcal{O}_{1}=2I\ ,$
$\mathrm{det}_{\mathcal{R}}\mathcal{O}_{2}=3I-W^{2}\ ,$
and
$\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+2}=2I\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+1}-(I+W^{2})\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q}\ .$ (53)
Proof.
The first two equalities for $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{1}$ and $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{2}$ are straightforward. For the one for ${\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+2}}$, expand the determinant along the top row. ∎
In the following, we seek to obtain a closed form for $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q}$. The recursive relation in (53) can be used to get
$\begin{pmatrix}\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+1}\\ \mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+2}\end{pmatrix}$ $=$ $\begin{pmatrix}0&I\\ -(I+W^{2})&2I\end{pmatrix}\begin{pmatrix}\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q}\\ \mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+1}\end{pmatrix}$ (54)
$=$ $\begin{pmatrix}0&I\\ -(I+W^{2})&2I\end{pmatrix}^{q}\begin{pmatrix}\mathrm{det}_{\mathcal{R}}\mathcal{O}_{1}\\ \mathrm{det}_{\mathcal{R}}\mathcal{O}_{2}\end{pmatrix}\ .$
Suppose $W$ is invertible in $\mathcal{R}$. The eigenvalues in $\mathcal{R}$ of $\Big{(}\begin{smallmatrix}0&I\\ -(I+W^{2})&2I\end{smallmatrix}\Big{)}$ are $I\pm iW$, and thus through diagonalization we can write
$\begin{pmatrix}0&I\\ -(I+W^{2})&2I\end{pmatrix}=\begin{pmatrix}I&I\\ I+iW&I-iW\end{pmatrix}\begin{pmatrix}I+iW&0\\ 0&I-iW\end{pmatrix}\\ \cdot\begin{pmatrix}\frac{1}{2}(I+iW^{-1})&-\frac{i}{2}W^{-1}\\ \frac{1}{2}(I-iW^{-1})&\phantom{-}\frac{i}{2}W^{-1}\end{pmatrix}\ ,$ (55)
with $\Big{(}\begin{smallmatrix}I&I\\ I+iW&I-iW\end{smallmatrix}\Big{)}$ and $\frac{1}{2}\Big{(}\begin{smallmatrix}I+iW^{-1}&-iW^{-1}\\ I-iW^{-1}&\phantom{-}iW^{-1}\end{smallmatrix}\Big{)}$ being inverses in $\mathcal{R}$. We use this to compute the $q$-th power of $\Big{(}\begin{smallmatrix}0&I\\ -(I+W^{2})&2I\end{smallmatrix}\Big{)}$, which equals
($\ (I+W^{2})\big{(}(I-iW)^{q}-(I+iW)^{q}\big{)}&(I+iW)^{q+1}-(I-iW)^{q+1}\end{pmatrix}\ ,$
and where the entries in the top row are not displayed since they are not needed for our purposes. Using this to compute the bottom row of (54), we get
$2iW\mathrm{det}_{\mathcal{R}}\mathcal{O}_{q+2}$ $=$ $\Big{[}(I+W^{2})\Big{(}(I-iW)^{q}-(I+iW)^{q}\Big{)}\Big{]}\mathrm{det}_{\mathcal{R}}\mathcal{O}_{1}$
$\qquad+\Big{[}(I+iW)^{q+1}-(I-iW)^{q+1}\Big{]}\mathrm{det}_{\mathcal{R}}\mathcal{O}_{2}$
$=$ $\Big{[}(I+W^{2})\Big{(}(I-iW)^{q}-(I+iW)^{q}\Big{)}\Big{]}2I$
$\qquad+\Big{[}(I+iW)^{q+1}-(I-iW)^{q+1}\Big{]}\Big{(}3I-W^{2}\Big{)}$
$=$ $(I-iW)^{q}\Big{[}(I+W^{2})2I-(I-iW)(3I-W^{2})\Big{]}$
$\ \ +\ (I+iW)^{q}\Big{[}-(I+W^{2})2I+(I+iW)(3I-W^{2})\Big{]}$
$=$ $(I-iW)^{q}\Big{[}-(I-iW)^{3}\Big{]}+(I+iW)^{q}\Big{[}(I+iW)^{3}\Big{]}$
$=$ $-(I-iW)^{q+3}+(I+iW)^{q+3}\ .$
Shift the values of $q$ by making $m-1=q+2$, and notice that
$(I+iW)^{m}-(I-iW)^{m}$ $=$ $\sum_{\ell=0}^{m}\binom{m}{\ell}I^{m-\ell}(iW)^{\ell}-\sum_{\ell=0}^{m}\binom{m}{\ell}I^{m-\ell}(-iW)^{\ell}$
$=$ $\sum_{\ell=0}^{m}\binom{m}{\ell}\big{(}i^{\ell}-(-i)^{\ell}\big{)}W^{\ell}$
$=$ $2iW\sum_{\ell=0}^{\lfloor\frac{m-1}{2}\rfloor}\binom{m}{2\ell+1}(-1)^{\ell}W^{2\ell}\ ,$
which is a polynomial in $W$. We now state the following.
Proposition B.2.
Let $m\geq 2$ and $W=G^{-1}H$. Then
$2iW\mathrm{det}_{\mathcal{R}}\mathcal{O}_{m-1}=(I+iW)^{m}-(I-iW)^{m}\ ,$
or equivalently
@@
Proof.
It only remains to justify the case in which $W$ is not invertible in $\mathcal{R}$. By the above arguments, for any invertible matrix, the value of the determinant of the corresponding block matrix will be in terms of the polynomial expression in (56), and hence the same expression will remain valid for $W$ by continuity and the density of the invertible matrices in $M_{d\times d}(\mathbb{C})$. ∎
@@
Consider $\gamma:B(0,1)\rightarrow\Gamma$ a parametrization of $\Gamma$ around $p$. The matrix $W$ is the matrix representation of the endomorphism $K_{p}$ with the basis for $\mathrm{D}_{p}\iota(\mathrm{T}_{p}\Gamma)$ given by $\left\{\mathrm{D}_{p}\iota\left(\frac{\partial}{\partial\gamma_{j}}\Big{\rvert}_{t}\right)\right\}_{j=1,\dots,d}$.
@@
Let $P_{m-1}$ be the polynomial defined by the right-hand side of (56), this is
$P_{m-1}(x)=\sum_{\ell=0}^{\lfloor\frac{m-1}{2}\rfloor}\binom{m}{2\ell+1}(-1)^{\ell}x^{2\ell}\ .$
By the spectral mapping theorem the eigenvalues of $P_{m-1}(W)$ are
@@
each with double the multiplicity as in $W$, together with the eigenvalue $P_{m-1}(0)=m$ with multiplicity $d-2r$ corresponding to the kernel of $W$. We present a first statement giving an expression for $\det\mathcal{O}_{m-1}$ in terms of the eigenvalues of $W$.
@@
$=$ $\prod_{\left\{\lambda\ :\ \lambda\ \mathrm{eigenvalue\ of}\ P_{m-1}(W)\right\}}\lambda$
@@
B.2 Some symplectic geometry
In this section we describe the behavior of $K_{p}$ when $\Gamma$ is isotropic or co-isotropic in order to express the eigenvalues ±$\pm i\lambda_{k}$ of $W$ needed to compute the determinant of $\mathcal{O}_{m-1}$.

### v2

@@ 5.2 Schatten norm estimates
being finite. For $p\geq 1$, $\left|\left|\cdot\right|\right|_{\scriptscriptstyle S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))}$ is a norm in $S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))$, while when $10<p<$ 1 it is a quasi-norm.
@@
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{\infty}_{0}(\Gamma)$. For any $p>0$,
@@ 5.3 An example
For $r\in(0,1)$ let $\Gamma=r\mathbb{S}^{1}=\{z\in\mathbb{C}:|z|=r\}$. If we use the parametrization $\gamma:(0,1)\rightarrow\Gamma$, $\gamma(\theta)=re^{2\pi i\theta}$, then $\mathrm{d}\sigma=2\pi\frac{r}{1-r^{2}}\mathrm{d}\theta$. Notice that in this case $d=1$ and $n=1$, thus $d^{\prime}=1$. Let $a\equiv 1$ in $\Gamma$ and denote $T_{r}:=T_{1\mathrm{d}\sigma}$ the associated Toeplitz operator. Then
@@ 5.3.1 Operator norm
from the known asymptotic formula $\mathrm{\mathbf{B}}\left(x,y\right)\sim\sqrt{2\pi}\frac{x^{x-\frac{1}{2}}y^{y-\frac{1}{2}}}{(x+y)^{x+y-\frac{1}{2}}}$ for large $x$ and $y$. Substituting this in the expression for $\lambda_{m^{*}}$ in (39) we get
@@ 5.3.2 Eigenvalue distribution comparison
Keeping with $\widehat{T}_{r}$ as defined above, given an interval $[t_{1},t_{2}]\subset\left(0,\frac{1}{(1-r^{2})^{2}}\right)$ Corollary 5.1 implies
@@
For fixed values of $r\in(0,1)$ and $t_{1}<t_{2}\leq(1-r^{2})^{-2}$ we have numerically counted the number of eigenvalues of $\widehat{T}_{r}$ in $[t_{1},t_{2}]$ for growing values of $\alpha$ using (39). Two comparisons between these numerical outputs and the expected results given by (40) and (41) are presented in the next two tables.
@@ B.1 The determinant as a block matrix
Recall from (23) that computing the determinant of the Hessian of $\varphi(t,\boldsymbol{0})$ relies on computing that of $\mathcal{O}_{m-1}$, which we have defined as the $(m-1)d\times(m-1)d$ matrix that expressed as a $(m-1)\times(m-1)$ block matrix with blocks of size $d\times d$ has the tridiagonal form
$\mathcal{O}_{m-1}=\begin{pmatrix}2I&-I-iW&{\scriptstyle\cdots}&0\\ -I+iW&2I&{\scriptstyle\cdots}&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&{\scriptstyle\cdots}&2I&-I-iW\\ 0&{\scriptstyle\cdots}&-I+iW&2I\end{pmatrix}.$
Here $W:=G^{-1}H$, where $G$ and $H$ are the matrices encoding the Riemannian metric $g_{p}$ and the skew-symmetric form $h_{p}$ on $\mathrm{T}_{p}\Gamma$ in terms of a parametrization $\gamma$ of $\Gamma$ with $\gamma(t)=p$. This same block matrix appeared in [11] in the Fock space situation.
Let $\mathcal{R}$ stand for the ring of $d\times d$ complex matrices generated by $I$ and $W$. Then $\mathcal{O}_{m-1}$ can also be regarded as a $(m-1)\times(m-1)$ matrix with entries in $\mathcal{R}$. As $\mathcal{R}$ is commutative, any such matrix possesses a well-defined determinant in $\mathcal{R}$, which we will denote $\mathrm{det}_{\mathcal{R}}$, and that possesses all the usual properties of determinants of matrices with complex entries. The relation between this determinant and the usual one $\det\mathcal{O}_{m-1}$ is given in Theorem 1 in [12]. From that theorem it follows that
@@
Notice that the right-hand side of the equation makes sense since $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{m-1}$ is an element of $\mathcal{R}$, and thus a $d\times d$ matrix with complex entries. This relation was used in [11] to compute $\det\mathcal{O}_{m-1}$. We cite the result therein expressing $\mathrm{det}_{\mathcal{R}}\mathcal{O}_{m-1}$ in terms of $W$, as all involved computations translate straightforwardly.
@@
Let $m\geq 2$. Then
@@
Consider $\gamma:B(0,1)\rightarrow\Gamma$ a parametrization of $\Gamma$ around $p$. Then $W$ is the matrix representation of the endomorphism $K_{p}$ with the basis for $\mathrm{D}_{p}\iota(\mathrm{T}_{p}\Gamma)$ given by $\left\{\mathrm{D}_{p}\iota\left(\frac{\partial}{\partial\gamma_{j}}\Big{\rvert}_{t}\right)\right\}_{j=1,\dots,d}$.
@@
Let $P_{m-1}$ be the polynomial defined by the right-hand side of (49). Then by the spectral mapping theorem the eigenvalues of $P_{m-1}(W)$ are
@@
each with double the multiplicity as in $W$, together with the eigenvalue $P_{m-1}(0)=m$ with multiplicity $d-2r$ corresponding to the kernel of $W$. We use this to express $\det\mathcal{O}_{m-1}$ in terms of the eigenvalues of $W$.
@@
B.2 The case of isotropic or co-isotropic manifolds
In this section we describe the behavior of $K_{p}$ when $\Gamma$ is isotropic or co-isotropic, in order to express the eigenvalues ±$\pm i\lambda_{k}$ of $W$ needed to compute the determinant of $\mathcal{O}_{m-1}$.